Lars:

What is the solution to the system of equations? y = x/2 + 2 y = x + 1

MusicGeek:

is that like x/2 like x divided by 2?

Lars:

x/2 is the fraction

MusicGeek:

okay. so u would set them equal to each other. so x/2+2=x+1. then ud multiply the whole thing by 2 to get rid of the fraction. now it would be x+4=2x+2. then subtract 1 x from both sides to get 4=x+2. then subtract 2 from both sides and u now have 2=x. so the answer is 2.

Lars:

Ohh okay that makes sense now, I see. thank you for that!

MusicGeek:

@lars wrote:
Ohh okay that makes sense now, I see. thank you for that!
you're welcome. lmk if u need anymore math help

Lars:

@musicgeek wrote:
@lars wrote:
Ohh okay that makes sense now, I see. thank you for that!
you're welcome. lmk if u need anymore math help
Okay I will I appreciate it